Operation Summary for January 4, 2010 - January 6, 2010 |
Monday Jan. 4 The day shift began with Operations monitoring the Tevatron (TeV) store 7477 (59.89E30) and the Recycler stash (164.9E10), with the Antiproton Source stacking (14.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations halted stacking from 2:06 to 2:44 PM to diagnose and repair a Pbar cooling problem. During the evening shift, Operations terminated store 7477 at 7:27 PM. Operators changed out a regulator for TeV power supply (T:HE15) and then performed orbit checks. Operations put the TeV into HEP shot setup, but the TeV quenched at 11:21 PM during the final proton load. Tuesday Jan. 5 The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash (456.8E10), with Pbar stacking (12.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators reported that the TeV low-level RF (LLRF) front end had caused the quench; diagnosis and repairs continued through to 6 AM. Operators performed orbit checks and then put the TeV into shot setup at 7:46 AM. During the day shift, the TeV shot setup was delayed when one transfer of antiprotons was lost; operators recovered and re-mined the Recycler. Operations established store 7484 at 9:38 AM with an average initial luminosity (AIL) of 267.53E30. At 1:38 PM, the TeV RF voltage drooped causing abort gap beam position problems. During the evening shift, Operations terminated store 7484 at 8:32 PM and put the TeV into shot setup. Operations established store 7485 at 10:05 PM with an AIL of 282.2E30. Wednesday Jan. 6 The midnight shift began with Operations monitoring store 7485 (191.93E30) and the stash (109.8E10), with Pbar stacking (14.0E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 6:20 AM, a TeV RF station (TRF4) tripped off. Operators will reset the station after the store ends. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 02/03/2010 email Fermilab |
| Security, Privacy, Legal | |